help me to solve this equations:(X+5)(X-5)=-7
4x-13/x=3/2
If you rest to a square of a number 16 units the result is 15 times the number.whats the number?

Answer by venugopalramana(3286)

SEE THE FOLLOWING EXAMPLE AND YOU CAN SOLVE YOUR PROBLEM.IN CASE OF DIFFICULTY COME BACK
I don't understanding any of this. can someone help??
(x-2)^2+x^2=10
Know how do I slove this???
1 solutions
Answer 11428 by venugopalramana(515)on 2005-12-20 10:24:17 
(X-2)^2+X^2=10..EXPANDING USING (A-B)^2 FORMULA
X^2-4X+4+X^2=10......TRANSPOSING AND SIMPLIFYING
2X^2-4X-6=0...DIVIDING BY 2 THROUGHOUT
X^2-2X-3=0
X^2-3X+X-3=0
X(X-3)+1(X-3)=0
(X-3)(X+1)=0
X-3 =0 .........................OR.................X+1=0
X=3.........OR..................X=-1